# Adaptive Data-Knowledge Alignment in Genetic Perturbation Prediction

### v1

@@ D.2 Reproducibility and Hyper-parameters
D.3 Full Results of Consistency Benchmark on Norman et al. (2019); Dixit et al. (2016); Adamson et al. (2016) Datasets
Table 2: Consistency benchmark in Figure 3,4
@@
Dataset Model ABL Stage Data Cons. Knowledge Cons. Balanced Cons.
@@
ALIGNED
baseline $0.5360_{\pm 0.0019}$ $0.3767_{\pm 0.0063}$ $0.4192_{\pm 0.0062}$
Align 1 $0.5040_{\pm 0.0106}$ $0.4307_{\pm 0.0197}$ $0.4575_{\pm 0.0125}$
Refine 1 $0.5252_{\pm 0.0042}$ $0.6549_{\pm 0.0044}$ $0.5697_{\pm 0.0024}$
Align 2 $0.5246_{\pm 0.0047}$ $0.6528_{\pm 0.0064}$ $0.5687_{\pm 0.0024}$
Refine 2 $0.5248_{\pm 0.0046}$ $0.6573_{\pm 0.0045}$ $0.5698_{\pm 0.0028}$
ALIGNED
baseline $0.5386_{\pm 0.0017}$ $0.3820_{\pm 0.0109}$ $0.4242_{\pm 0.0102}$
Align 1 $0.5154_{\pm 0.0163}$ $0.4133_{\pm 0.0366}$ $0.4447_{\pm 0.0260}$
Refine 1 $0.5270_{\pm 0.0023}$ $0.6555_{\pm 0.0021}$ $0.5714_{\pm 0.0022}$
Align 2 $0.5261_{\pm 0.0032}$ $0.6559_{\pm 0.0024}$ $0.5711_{\pm 0.0025}$
Refine 2 $0.5261_{\pm 0.0033}$ $0.6596_{\pm 0.0019}$ $0.5717_{\pm 0.0029}$
@@
ALIGNED
baseline $0.4207_{\pm 0.0028}$ $0.3930_{\pm 0.0184}$ $0.4052_{\pm 0.0113}$
Align 1 $0.4126_{\pm 0.0022}$ $0.4245_{\pm 0.0168}$ $0.4173_{\pm 0.0069}$
Refine 1 $0.4188_{\pm 0.0039}$ $0.6591_{\pm 0.0036}$ $0.4718_{\pm 0.0041}$
Align 2 $0.4181_{\pm 0.0036}$ $0.6579_{\pm 0.0042}$ $0.4711_{\pm 0.0038}$
Refine 2 $0.4180_{\pm 0.0036}$ $0.6630_{\pm 0.0014}$ $0.4711_{\pm 0.0037}$
ALIGNED
baseline $0.4509_{\pm 0.0068}$ $0.3808_{\pm 0.0033}$ $0.4069_{\pm 0.0040}$
Align 1 $0.4534_{\pm 0.0102}$ $0.4147_{\pm 0.0251}$ $0.4289_{\pm 0.0126}$
Refine 1 $0.4585_{\pm 0.0048}$ $0.6367_{\pm 0.0122}$ $0.5082_{\pm 0.0046}$
Align 2 $0.4588_{\pm 0.0045}$ $0.6376_{\pm 0.0130}$ $0.5082_{\pm 0.0046}$
Refine 2 $0.4589_{\pm 0.0047}$ $0.6555_{\pm 0.0023}$ $0.5110_{\pm 0.0047}$
@@
ALIGNED
baseline $0.2888_{\pm 0.0030}$ $0.3554_{\pm 0.0293}$ $0.3128_{\pm 0.0077}$
Align 1 $0.4767_{\pm 0.0108}$ $0.3725_{\pm 0.0315}$ $0.4063_{\pm 0.0281}$
Refine 1 $0.4682_{\pm 0.0102}$ $0.4111_{\pm 0.0261}$ $0.4325_{\pm 0.0179}$
Align 2 $0.4672_{\pm 0.0086}$ $0.6467_{\pm 0.0056}$ $0.5178_{\pm 0.0078}$
Refine 2 $0.4674_{\pm 0.0078}$ $0.6418_{\pm 0.0083}$ $0.5178_{\pm 0.0077}$
ALIGNED
baseline $0.5436_{\pm 0.0059}$ $0.4319_{\pm 0.0317}$ $0.4692_{\pm 0.0268}$
Align 1 $0.4925_{\pm 0.0450}$ $0.5589_{\pm 0.1079}$ $0.5023_{\pm 0.0226}$
Refine 1 $0.5225_{\pm 0.0177}$ $0.6321_{\pm 0.0097}$ $0.5641_{\pm 0.0142}$
Align 2 $0.5205_{\pm 0.0168}$ $0.6328_{\pm 0.0105}$ $0.5611_{\pm 0.0127}$
Refine 2 $0.5235_{\pm 0.0179}$ $04_{\pm 0.0026} 0.5665±0$$65_{\pm 0.0143}$

### v2

@@ F.3 Interaction Recovery
F.6 Full Results of Consistency Benchmark on Norman et al. (2019); Dixit et al. (2016); Adamson et al. (2016) Datasets
Evaluations of Section 4.1 on train and test data in precision and recall are shown in Table 11. Full benchmark results of Section 4.1 are shown in Table 12.
Table 11: Precision and recall in Figure 3,5. “A” is an alignment stage, and “R” is a refinement stage. ALIGNED with “A-R” executes alignment and then refinement.
@@
Dataset Model ALIGNED Stage Train Data Test Data Knowledge Base
Precision Recall Precision Recall Precision Recall
Norman MLP MLP only $0.556$ $0.544$ $0.543$ $0.541$ $0.568$ $0.427$
A $0.510$ $0.539$ $0.503$ $0.537$ $0.593$ $0.465$
A-R $0.553$ $0.535$ $0.537$ $0.534$ $0.644$ $0.652$
GNN GNN only $0.552$ $0.556$ $0.552$ $0.553$ $0.576$ $0.433$
A $0.524$ $0.544$ $0.523$ $0.543$ $0.597$ $0.458$
A-R $0.553$ $0.538$ $0.545$ $0.540$ $0.653$ $0.655$
Dixit MLP MLP only $0.498$ $0.449$ $0.494$ $0.448$ $0.708$ $0.476$
A $0.474$ $0.445$ $0.454$ $0.445$ $0.599$ $0.483$
A-R $0.503$ $0.446$ $0.462$ $0.441$ $0.590$ $0.645$
GNN GNN only $0.493$ $0.489$ $0.514$ $0.505$ $0.604$ $0.502$
A $0.480$ $0.499$ $0.480$ $0.504$ $0.563$ $0.525$
A-R $0.485$ $0.498$ $0.484$ $0.499$ $0.616$ $0.638$
Adamson MLP MLP only $0.559$ $0.463$ $0.524$ $0.467$ $0.587$ $0.448$
A $0.537$ $0.464$ $0.486$ $0.467$ $0.533$ $0.460$
A-R $0.571$ $0.461$ $0.496$ $0.457$ $0.562$ $0.643$
GNN GNN only $0.561$ $0.555$ $0.578$ $0.579$ $0.577$ $0.502$
A $0.511$ $0.528$ $0.531$ $0.566$ $0.676$ $0.634$
A-R $0.557$ $0.543$ $0.553$ $0.560$ $0.633$ $0.645$
Table 12: Consistency benchmark in Figure 3,5. “A” is an alignment stage, and “R” is a refinement stage. ALIGNED with “A-R” executes alignment and then refinement.
Dataset Model ALIGNED Data Knowledge Balanced
Stages Cons. Cons. Cons.
@@
ALIGNED (MLP) MLP only $0.5360_{\pm 0.0019}$ $0.3767_{\pm 0.0063}$ $0.4192_{\pm 0.0062}$
A $0.5040_{\pm 0.0106}$ $0.4307_{\pm 0.0197}$ $0.4575_{\pm 0.0125}$
A-R $0.5252_{\pm 0.0042}$ $0.6549_{\pm 0.0044}$ $0.5697_{\pm 0.0024}$
A-R-A $0.5246_{\pm 0.0047}$ $0.6528_{\pm 0.0064}$ $0.5687_{\pm 0.0024}$
A-R-A-R $0.5248_{\pm 0.0046}$ $0.6573_{\pm 0.0045}$ $0.5698_{\pm 0.0028}$
ALIGNED (GNN) GNN only $0.5386_{\pm 0.0017}$ $0.3820_{\pm 0.0109}$ $0.4242_{\pm 0.0102}$
A $0.5154_{\pm 0.0163}$ $0.4133_{\pm 0.0366}$ $0.4447_{\pm 0.0260}$
A-R $0.5270_{\pm 0.0023}$ $0.6555_{\pm 0.0021}$ $0.5714_{\pm 0.0022}$
A-R-A $0.5261_{\pm 0.0032}$ $0.6559_{\pm 0.0024}$ $0.5711_{\pm 0.0025}$
A-R-A-R $0.5261_{\pm 0.0033}$ $0.6596_{\pm 0.0019}$ $0.5717_{\pm 0.0029}$
@@
ALIGNED (MLP) MLP only $0.4207_{\pm 0.0028}$ $0.3930_{\pm 0.0184}$ $0.4052_{\pm 0.0113}$
A $0.4126_{\pm 0.0022}$ $0.4245_{\pm 0.0168}$ $0.4173_{\pm 0.0069}$
A-R $0.4188_{\pm 0.0039}$ $0.6591_{\pm 0.0036}$ $0.4718_{\pm 0.0041}$
A-R-A $0.4181_{\pm 0.0036}$ $0.6579_{\pm 0.0042}$ $0.4711_{\pm 0.0038}$
A-R-A-R $0.4180_{\pm 0.0036}$ $0.6630_{\pm 0.0014}$ $0.4711_{\pm 0.0037}$
ALIGNED (GNN) GNN only $0.4509_{\pm 0.0068}$ $0.3808_{\pm 0.0033}$ $0.4069_{\pm 0.0040}$
A $0.4534_{\pm 0.0102}$ $0.4147_{\pm 0.0251}$ $0.4289_{\pm 0.0126}$
A-R $0.4585_{\pm 0.0048}$ $0.6367_{\pm 0.0122}$ $0.5082_{\pm 0.0046}$
A-R-A $0.4588_{\pm 0.0045}$ $0.6376_{\pm 0.0130}$ $0.5082_{\pm 0.0046}$
A-R-A-R $0.4589_{\pm 0.0047}$ $0.6555_{\pm 0.0023}$ $0.5110_{\pm 0.0047}$
@@
ALIGNED (MLP) MLP only $0.2888_{\pm 0.0030}$ $0.3554_{\pm 0.0293}$ $0.3128_{\pm 0.0077}$
A $0.4767_{\pm 0.0108}$ $0.3725_{\pm 0.0315}$ $0.4063_{\pm 0.0281}$
A-R $0.4682_{\pm 0.0102}$ $0.4111_{\pm 0.0261}$ $0.4325_{\pm 0.0179}$
A-R-A $0.4672_{\pm 0.0086}$ $0.6467_{\pm 0.0056}$ $0.5178_{\pm 0.0078}$
A-R-A-R $0.4674_{\pm 0.0078}$ $0.6418_{\pm 0.0083}$ $0.5178_{\pm 0.0077}$
ALIGNED (GNN) GNN only $0.5436_{\pm 0.0059}$ $0.4319_{\pm 0.0317}$ $0.4692_{\pm 0.0268}$
A $0.4925_{\pm 0.0450}$ $0.5589_{\pm 0.1079}$ $0.5023_{\pm 0.0226}$
A-R $0.5225_{\pm 0.0177}$ $0.6321_{\pm 0.0097}$ $0.5641_{\pm 0.0142}$
A-R-A $0.5205_{\pm 0.0168}$ $0.6328_{\pm 0.0105}$ $0.5611_{\pm 0.0127}$
A-R-A-R $0.5235_{\pm 0.0179}$ $04_{\pm 0.0026} 0.5665±0$$65_{\pm 0.0143}$